OpenStudy (anonymous):

Find the equation of the curve that joins the following points. Hint: Use common differences. (4 marks) a) The zeros of the function are -3, 1 and 5. The y intercept is 15. Find the function. b) (1,1), (0,1), (2,-5), (-1,-5)

OpenStudy (anonymous):

@ganeshie8 @ParthKohli hey could you guys please help me i have no clue how to do this.

OpenStudy (anonymous):

@mathstudent55 @AndrewTheCookie @mathmate @TheSmartOne

OpenStudy (tkhunny):

a or b? The question seems to refer to b. Show Your Work, please.

OpenStudy (anonymous):

ok 1 sec i scan it

OpenStudy (anonymous):

i got the answer for a) but i did not use common difference, dont know how to use it in this situation, and i dont know how to do part b)

OpenStudy (tkhunny):

b) , (-1,-5),(0,1),(1,1), (2,-5) From -1 to 0 1 - (-5) = 6 From 0 to 1 1 - 1 = 0 From 1 to 2 -5 - 1 = -6

OpenStudy (anonymous):

dont you have to rearrange it ? then subtrac till you get a common difference ?

OpenStudy (anonymous):

im not sure if i should rearrange it or not..

OpenStudy (tkhunny):

Yes. I rearranged and showed the first difference.

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

let me try it out then thanks. I try it then get bak to you :)

OpenStudy (anonymous):

hey could you tell me if a) is correct @tkhunny

OpenStudy (cwrw238):

yes a is correct

OpenStudy (anonymous):

thank you :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

a rough diagram should help with part b

OpenStudy (cwrw238):